OpenStudy (anonymous):

What is the y-intercept of the equation? y = 0.5(4)x

OpenStudy (jdoe0001):

to get the y-intercept, set x=0, and solve for "y"

OpenStudy (anonymous):

so y=0.5(4)0?

OpenStudy (anonymous):

and how so i solve for y

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

is it 0.5 ? the answer ?

OpenStudy (jdoe0001):

well, you isolate "y" on either side, left or right

OpenStudy (anonymous):

yeah i remebered :) is the answer 0.5 ?

OpenStudy (jdoe0001):

let's see \( \bf y = 0.5(4)x \implies y = 0.5\times 4\times x\\ y =0.5 \times 4\times 0\\ y = \square?\)